OpenStudy (anonymous):

From a pack of 52 cards, 3 cards are selected at random. What is the probability of getting more number of kings than queens?

OpenStudy (anonymous):

52C3=52*51*50/2*3=22100

OpenStudy (anonymous):

4C2*4c1*44c0/52 C3=6*4/22100

OpenStudy (anonymous):

or else (4c1*4c0*44c1+4c2*4c1*44c0+4c3*44c0*4c0)/52c3

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

or else 4/52*3/51*4/50+4/52*3/51*2/50

ganeshie8 (ganeshie8):

lets see the possible cases first

ganeshie8 (ganeshie8):

case 1 : `1 king`, 2 others case 2a : `2 king`, 1 others case 2b : `2 king`, `1 queen` case 3 : `3 kings`

ganeshie8 (ganeshie8):

In all the above cases, kings are more than queens, yes ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(48c2*4c1+4c2*48c1+4c2*4c1+4c3)/52c3

ganeshie8 (ganeshie8):

careful, in the first case you should not include queens

ganeshie8 (ganeshie8):

kings = 4 queens = 4 others = 44 ----------------------- total = 52

OpenStudy (anonymous):

okay

ganeshie8 (ganeshie8):

case 1 : `1 king`, 2 others : \(4C1*44C2\) case 2a : `2 king`, 1 others : \(4C2*44C1\) case 2b : `2 king`, `1 queen` : \(4C2*4C1\) case 3 : `3 kings` : \(4C3\)

ganeshie8 (ganeshie8):

add them all up to get the number of favorable outcomes

OpenStudy (anonymous):

4076/22100

ganeshie8 (ganeshie8):

Correct !

OpenStudy (anonymous):

(4c1*4c0*44c1+4c2*4c1*44c0+4c3*44c0*4c0)/52c3

ganeshie8 (ganeshie8):

you mean this : ( `4c1*4c0*44c2` + `4c2*4c0*44c1` + `4c2*4c1*44c0` + `4c3*44c0*4c0`)/52c3 right ?

OpenStudy (anonymous):

4096/5203,4076/5203,1024/5203,2048/5203,36/5203

OpenStudy (anonymous):

options are:4096/5203,4076/5203,1024/5203,2048/5203,36/5203

ganeshie8 (ganeshie8):

your options are wrong, i worked this question few months back... let me see if i can find the question with correct options

OpenStudy (anonymous):

i think wrong options are given

OpenStudy (anonymous):

okay.3q .u r right.:-)

OpenStudy (anonymous):

i am so confused with options so tried to get denominator 5203.3q

OpenStudy (anonymous):

i am clear now

ganeshie8 (ganeshie8):

ikr... i had spent 1 fully day on and off on this problem before realizing that the options were wrong by googling haha

OpenStudy (anonymous):

haha...........